# Stability of Poiseuille Flow of Navier-Stokes Equations on $\mathbb{R}^{2}$

### v1

@@ Abstract
We consider solutions to the Navier–Stokes equations on $\mathbb{R}^{2}$ close to the Poiseuille flow with viscosity $10<\nu<$ 1. For the linearized problem, we prove that when the $x$-frequency satisfy $|k|\geq\nu^{-\frac{1}{3}}$, the perturbation decays on a time-scale proportional to $\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}$. Since it decays faster than the heat equation, this phenomenon is referred to as enhanced dissipation. Then we concern the non-linear equations. We show that if the initial perturbation $\omega_{in}$ is of size at most $\nu^{\frac{7}{12}}$ in an anisotropic Sobolev space, then it remains so for all time.
@@ 1 Introduction
$U^{S}(x,y)=(y^{2},0),P^{S}(x,y)=2\nu x.$
This flow is important and one reason is that it is the simplest non-trivial example of a shear flow on $\mathbb{R}^{2}$ besides the Couette flow $U=(y,0)$.
@@
For the linear problem, we prove that when the $x$-frequency satisfy $|k|\geq\nu^{-\frac{1}{3}}$, the perturbation decays on a time-scale proportional to $\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}$. In other words, enhanced dissipation exists. We will see this in detail in Section 2. For the non-linear problem, we prove a quantitative stability threshold for the initial perturbation $\omega_{in}$. Namely, we show that for suitable anisotropic Sobolev norms $\|\cdot\|_{X}$ and $\|\cdot\|_{Y}$, that if $\|\omega_{in}\|_{X}\lesssim\nu^{\frac{7}{12}}$, then the corresponding solution $\omega$ of (1.3) satisfies $\|\omega\|_{Y}\ll 1$ for all $t>0$ and $\|\omega\|_{Y}\to 0$ as $t\to\infty$. Before presenting the main results, we first introduce the following notation.
@@
We denote the corresponding Hamiltonian and Laplacian by $\nabla_{k}\coloneqq(ik,\partial_{y})$ and $\Delta_{k}\coloneqq\partial_{y}^{2}-|k|^{2}$.
@@
$+\|\nu^{-\frac{1}{3}}\langle\partial_{x}\rangle^{m}\langle\nu^{\frac{1}{3}}\partial_{x}\rangle^{\frac{1}{4}}\partial_{y}\Delta^{-1}\omega_{in}\|_{L_{x,y}^{2}}+\|y\omega_{in,k}\|_{L_{k}^{\infty}L_{y}^{2}}+\|\nabla_{k}\Delta_{k}^{-1}\omega_{in,k}\|_{L_{k}^{\infty}L_{y}^{2}}\leq\delta\nu^{\frac{7}{12}},$
@@
$+$ $\|\nu^{-\frac{1}{3}}\langle c\lambda_{\nu}^{pl}(\partial_{x})t\rangle^{J}\langle\partial_{x}\rangle^{m}\langle\nu^{\frac{1}{3}}\partial_{x}\rangle^{\frac{1}{4}}\partial_{y}\Delta^{-1}\omega\|_{L_{x,y}^{2}}+\|y\omega_{k}\|_{L_{k}^{\infty}L_{y}^{2}}+\|\nabla_{k}\Delta_{k}^{-1}\omega_{k}\|_{L_{k}^{\infty}L_{y}^{2}}\leq\epsilon.$
@@
$\lambda_{\nu}^{pl}(k)=\begin{cases*}\nu^{\frac{1}{2}}|k|^{\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu|k|^{2},|k|\leq\nu^{-\frac{1}{3}}.\end{cases*}$
@@ 2 Linearized Problem
$E_{k}=\frac{1}{2}\|\omega_{k}\|_{2}^{2}+\frac{1}{2}c_{\alpha}\alpha_{k}\|\nabla_{k}\omega_{k}\|_{2}^{2}+2c_{\beta}\beta_{k}\mathrm{Re}\langle iky\omega_{k}\partial_{y}\omega_{k}\rangle+\frac{1}{2}c_{\gamma}\gamma_{k}(\|y\omega_{k}\|_{2}^{2}+2\|\nabla_{k}\psi_{k}\|_{2}^{2}),$
@@
$\alpha_{k}=\begin{cases*}\nu^{\frac{1}{2}}|k|^{-\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu^{\frac{2}{3}},|k|\leq\nu^{-\frac{1}{3}},\end{cases*}\beta_{k}=\begin{cases*}|k|^{-1},|k|\geq\nu^{-\frac{1}{3}},\\ \nu^{\frac{1}{3}},|k|\leq\nu^{-\frac{1}{3}},\end{cases*}\gamma_{k}=\begin{cases*}\nu^{-\frac{1}{2}}|k|^{\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu^{-\frac{2}{3}},|k|\leq\nu^{-\frac{1}{3}},\end{cases*}\lambda_{k}=\begin{cases*}\nu^{\frac{1}{2}}|k|^{\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu|k|^{2},|k|\leq\nu^{-\frac{1}{3}},\end{cases*}$
@@
$\frac{\mathrm{d}E_{k}}{\mathrm{d}t}\leq-4cD_{k}-4c\lambda_{k}E_{k}.$ (2.4)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{k}\|_{2}^{2}=2\mathrm{Re}\int(iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\overline{\omega_{k}}\mathrm{d}y=-2\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}.$
@@
$=|k|^{2}(-2\nu\|\nabla_{k}\omega_{k}\|_{2}^{2})+2\mathrm{Re}\int\partial_{y}(iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\partial_{y}\overline{\omega_{k}}\mathrm{d}y$
@@
Using $\|\Delta_{k}\omega_{k}\|_{2}^{2}=\|\nabla_{k}^{2}\omega_{k}\|_{2}^{2}=\|k\nabla_{k}\omega_{k}\|_{2}^{2}+\|\partial_{y}\nabla_{k}\omega_{k}\|_{2}^{2}$ again, and then (2.6) follows. For (2.7), we directly compute by (2.2),
@@
$=$ $\nu\mathrm{Re}\int iky\Delta_{k}\omega_{k}\partial_{y}\overline{\omega_{k}}+iky\omega_{k}\partial_{y}\Delta_{k}\overline{\omega_{k}}\mathrm{d}y$
$+\mathrm{Re}\int k^{2}y^{3}\omega_{k}\partial_{y}\overline{\omega_{k}}-k^{2}y\omega_{k}\partial_{y}(y^{2}\overline{\omega_{k}})\mathrm{d}y$
$+\mathrm{Re}\int-2k^{2}y\psi_{k}\partial_{y}\overline{\omega_{k}}+2k^{2}y\omega_{k}\partial_{y}\overline{\psi_{k}}\mathrm{d}y.$
@@
$\nu\mathrm{Re}\int iky\Delta_{k}\omega_{k}\partial_{y}\overline{\omega_{k}}+iky\omega_{k}\partial_{y}\Delta_{k}\overline{\omega_{k}}\mathrm{d}y=-2\nu\mathrm{Re}\langle\Delta_{k}\omega_{k},iky\partial_{y}\omega_{k}\rangle,$
@@
$\mathrm{Re}\int k^{2}y^{3}\omega_{k}\partial_{y}\overline{\omega_{k}}-k^{2}y\omega_{k}\partial_{y}(y^{2}\overline{\omega_{k}})\mathrm{d}y=-2|k|^{2}\|y\omega_{k}\|_{2}^{2}.$
@@
$\mathrm{Re}\int-2k^{2}y\psi_{k}\partial_{y}\overline{\omega_{k}}+2k^{2}y\omega_{k}\partial_{y}\overline{\psi_{k}}\mathrm{d}y=-4|k|^{2}\|\partial_{y}\psi_{k}\|_{2}^{2},$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|y\omega_{k}\|_{2}^{2}$ $=2\mathrm{Re}\int y(-iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\overline{y\omega_{k}}\mathrm{d}y$
$=4\mathrm{Re}\int iky^{2}\psi_{k}\Delta_{k}\overline{\psi_{k}}\mathrm{d}y+2\nu\mathrm{Re}\int y^{2}\Delta_{k}\omega_{k}\overline{\omega_{k}}\mathrm{d}y$
$=-8\mathrm{Re}\langle iky\psi_{k},\partial_{y}\psi_{k}\rangle-4\nu\mathrm{Re}\int y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y-2\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}.$
@@
$\int y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y=-\int\partial_{y}(y\overline{\omega_{k}})\omega_{k}\mathrm{d}y=-\int(\overline{\omega_{k}}+y\partial_{y}\overline{\omega_{k}})\omega_{k}\mathrm{d}y=-\|\omega_{k}\|_{2}^{2}-\int y\partial_{y}\overline{\omega_{k}}\omega_{k}\mathrm{d}y,$
@@
$\mathrm{Re}\int y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y=\frac{1}{2}\left(\int y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y+\int y\partial_{y}\overline{\omega_{k}}\omega_{k}\mathrm{d}y\right)=-\frac{1}{2}\|\omega_{k}\|_{2}^{2}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla_{k}\psi_{k}\|_{2}^{2}=-2\mathrm{Re}\int\partial_{t}\Delta_{k}\psi_{k}\overline{\psi_{k}}\mathrm{d}y=-2\mathrm{Re}\int(-iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\overline{\psi_{k}}\mathrm{d}y,$
@@
$2\mathrm{Re}\int iky^{2}\Delta_{k}\psi_{k}\overline{\psi_{k}}\mathrm{d}y=2\mathrm{Re}\int-\partial_{y}(iky^{2}\overline{\psi_{k}})\partial_{y}\psi_{k}-ik^{3}y^{2}\psi_{k}\overline{\psi_{k}}\mathrm{d}y=4\mathrm{Re}\langle iky\psi_{k},\partial_{y}\psi_{k}\rangle.\qed$
@@
$\frac{\mathrm{d}E_{k}}{\mathrm{d}t}=$ $-\big{(}c_{\gamma}\gamma_{k}\nu\|\omega_{k}\|_{2}^{2}+\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}+c_{\alpha}\alpha_{k}\nu\|\Delta_{k}\omega_{k}\|_{2}^{2}+4c_{\beta}\beta_{k}|k|^{2}\|y\omega_{k}\|_{2}^{2}+c_{\gamma}\gamma_{k}\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}$
$+8c_{\beta}\beta_{k}|k|^{2}\|\partial_{y}\psi_{k}\|_{2}^{2}\big{)}-2c_{\alpha}\alpha_{k}\mathrm{Re}\langle iky\omega_{k},\partial_{y}\omega_{k}\rangle-4c_{\beta}\beta_{k}\nu\mathrm{Re}\langle\Delta_{k}\omega_{k},iky\partial_{y}\omega_{k}\rangle.$
@@
$\lambda_{k}E_{k}$ $\leq\nu\|k\omega_{k}\|_{2}^{2}+\alpha_{k}\nu\|k\nabla_{k}\omega_{k}\|_{2}^{2}+\gamma_{k}\nu|k|^{2}\|y\omega_{k}\|_{2}^{2}+\gamma_{k}\nu\|k\partial_{y}\psi_{k}\|_{2}^{2}$
$\lesssim\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}+\alpha_{k}\nu\|\Delta_{k}\omega_{k}\|_{2}^{2}+\nu^{\frac{1}{3}}|k|^{2}\|y\omega_{k}\|_{2}^{2}+\gamma_{k}\nu\|\omega_{k}\|_{2}^{2}\lesssim I.\qed$

### v4

@@ Abstract
We consider solutions to the Navier–Stokes equations on $\mathbb{R}^{2}$ close to the Poiseuille flow with viscosity $10<\nu<$ 1. For the linearized problem, we prove that when the $x$-frequency satisfies $|k|\geq\nu^{-\frac{1}{3}}$, the perturbation decays on a time-scale proportional to $\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}$. Since it decays faster than the heat equation, this phenomenon is referred to as enhanced dissipation. Then we concern the non-linear equations. We show that if the initial perturbation $\omega_{in}$ is at most of size $\nu^{\frac{7}{3}}$ in an anisotropic Sobolev space, then the size of the perturbation remains no more than twice the size of its initial value.
Keywords:Navier–Stokes equations, Poiseuille flow, stability
@@ 1 Introduction
$U^{P}(x,y)=(y^{2},0),P^{P}(x,y)=2\nu x.$
This flow is important and one reason is that it is the simplest non-trivial example of a shear flow on $\mathbb{R}^{2}$ besides the Couette flow $U^{C}=(y,0)$.
@@
For the linear problem, we prove that when the $x$-frequency satisfies $|k|\geq\nu^{-\frac{1}{3}}$, the perturbation decays on a time-scale proportional to $\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}$. In other words, enhanced dissipation exists. We will see this in detail in Section 2. For the non-linear problem, we prove a quantitative stability threshold for the initial perturbation $\omega_{in}$. Namely, we show that for suitable anisotropic Sobolev norms $\|\cdot\|_{X}$ and $\|\cdot\|_{Y}$, that if $\|\omega_{in}\|_{X}\lesssim\nu^{\frac{7}{3}}$, then the corresponding solution $\omega$ of (1.3) satisfies $\|\omega\|_{Y}\ll 1$ for all $t>0$ and $\|\omega\|_{Y}\to 0$ as $t\to\infty$. Before presenting the main results, we first introduce the following notations.
@@
We denote the corresponding Hamiltonian and Laplacian by $\nabla_{k}\coloneqq(ik,\partial_{y})$ and $\Delta_{k}\coloneqq\partial_{y}^{2}-|k|^{2}$. We use the standard notation $\langle\cdot\rangle\coloneqq\sqrt{1+\cdot^{2}}$. Given two quantities $A,B\geq 0$, we write $A\lesssim B$ to indicate that there exists a constant $C>0$ such that $A\leq CB$. If $A\lesssim B$ and $B\lesssim A$, we write $A\approx B$.
@@
$+\|\nu^{-\frac{1}{3}}\langle\partial_{x}\rangle^{m}\langle\nu^{\frac{1}{3}}\partial_{x}\rangle^{\frac{1}{4}}\partial_{y}\Delta^{-1}\omega_{in}\|_{L_{x,y}^{2}}+\|y\omega_{in,k}\|_{L_{k}^{\infty}L_{y}^{2}}+\|\nabla_{k}\Delta_{k}^{-1}\omega_{in,k}\|_{L_{k}^{\infty}L_{y}^{2}}\leq\delta\nu^{\frac{7}{3}},$
@@
$+\|\nu^{-\frac{1}{3}}\langle c\lambda_{\nu}^{pl}(\partial_{x})t\rangle^{J}\langle\partial_{x}\rangle^{m}\langle\nu^{\frac{1}{3}}\partial_{x}\rangle^{\frac{1}{4}}\partial_{y}\Delta^{-1}\omega\|_{L_{x,y}^{2}}+\|y\omega_{k}\|_{L_{k}^{\infty}L_{y}^{2}}+\|\nabla_{k}\Delta_{k}^{-1}\omega_{k}\|_{L_{k}^{\infty}L_{y}^{2}}\leq 2\epsilon.$
@@
$\lambda_{\nu}^{pl}(k)=\begin{cases*}\nu^{\frac{1}{2}}|k|^{\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu|k|^{2},|k|<\nu^{-\frac{1}{3}}.\end{cases*}$
@@ 2 Linearized Problem
$E_{k}=\frac{1}{2}\|\omega_{k}\|_{2}^{2}+\frac{1}{2}c_{\alpha}\alpha_{k}\|\nabla_{k}\omega_{k}\|_{2}^{2}+2c_{\beta}\beta_{k}\mathrm{Re}\langle iky\omega_{k},\partial_{y}\omega_{k}\rangle+\frac{1}{2}c_{\gamma}\gamma_{k}(\|y\omega_{k}\|_{2}^{2}+2\|\nabla_{k}\psi_{k}\|_{2}^{2}),$
@@
$\alpha_{k}=\begin{cases*}\nu^{\frac{1}{2}}|k|^{-\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu^{\frac{2}{3}},|k|<\nu^{-\frac{1}{3}},\end{cases*}\beta_{k}=\begin{cases*}|k|^{-1},|k|\geq\nu^{-\frac{1}{3}},\\ \nu^{\frac{1}{3}},|k|<\nu^{-\frac{1}{3}},\end{cases*}\gamma_{k}=\begin{cases*}\nu^{-\frac{1}{2}}|k|^{\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu^{-\frac{2}{3}},|k|<\nu^{-\frac{1}{3}},\end{cases*}$
and constants $c_{\alpha},c_{\beta},c_{\gamma}$ satisfy
@@
$\frac{\mathrm{d}E_{k}}{\mathrm{d}t}\leq-4cD_{k}-4c\lambda_{k}E_{k},$ (2.4)
where
$\lambda_{k}=\begin{cases*}\nu^{\frac{1}{2}}|k|^{\frac{1}{2}},|k|\geq\nu^{-\frac{1}{3}},\\ \nu|k|^{2},|k|<\nu^{-\frac{1}{3}}.\end{cases*}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{k}\|_{2}^{2}=2\mathrm{Re}\int_{\mathbb{R}}(iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\overline{\omega_{k}}\mathrm{d}y=-2\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}.$
@@
$=|k|^{2}(-2\nu\|\nabla_{k}\omega_{k}\|_{2}^{2})+2\mathrm{Re}\int_{\mathbb{R}}\partial_{y}(iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\partial_{y}\overline{\omega_{k}}\mathrm{d}y$
@@
Using $\|\Delta_{k}\omega_{k}\|_{2}^{2}=\|\nabla_{k}^{2}\omega_{k}\|_{2}^{2}=\|k\nabla_{k}\omega_{k}\|_{2}^{2}+\|\partial_{y}\nabla_{k}\omega_{k}\|_{2}^{2}$, and then (2.6) follows. For (2.7), we directly compute by (2.2),
@@
$=$ $\nu\mathrm{Re}\int_{\mathbb{R}}iky\Delta_{k}\omega_{k}\partial_{y}\overline{\omega_{k}}+iky\omega_{k}\partial_{y}\Delta_{k}\overline{\omega_{k}}\mathrm{d}y$
$+\mathrm{Re}\int_{\mathbb{R}}k^{2}y^{3}\omega_{k}\partial_{y}\overline{\omega_{k}}-k^{2}y\omega_{k}\partial_{y}(y^{2}\overline{\omega_{k}})\mathrm{d}y$
$+\mathrm{Re}\int_{\mathbb{R}}-2k^{2}y\psi_{k}\partial_{y}\overline{\omega_{k}}+2k^{2}y\omega_{k}\partial_{y}\overline{\psi_{k}}\mathrm{d}y.$
@@
$\nu\mathrm{Re}\int_{\mathbb{R}}iky\Delta_{k}\omega_{k}\partial_{y}\overline{\omega_{k}}+iky\omega_{k}\partial_{y}\Delta_{k}\overline{\omega_{k}}\mathrm{d}y=-2\nu\mathrm{Re}\langle\Delta_{k}\omega_{k},iky\partial_{y}\omega_{k}\rangle,$
@@
$\mathrm{Re}\int_{\mathbb{R}}k^{2}y^{3}\omega_{k}\partial_{y}\overline{\omega_{k}}-k^{2}y\omega_{k}\partial_{y}(y^{2}\overline{\omega_{k}})\mathrm{d}y=-2|k|^{2}\|y\omega_{k}\|_{2}^{2}.$
@@
$\mathrm{Re}\int_{\mathbb{R}}-2k^{2}y\psi_{k}\partial_{y}\overline{\omega_{k}}+2k^{2}y\omega_{k}\partial_{y}\overline{\psi_{k}}\mathrm{d}y=-4|k|^{2}\|\partial_{y}\psi_{k}\|_{2}^{2},$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|y\omega_{k}\|_{2}^{2}$ $=2\mathrm{Re}\int_{\mathbb{R}}y(-iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\overline{y\omega_{k}}\mathrm{d}y$
$=4\mathrm{Re}\int_{\mathbb{R}}iky^{2}\psi_{k}\Delta_{k}\overline{\psi_{k}}\mathrm{d}y+2\nu\mathrm{Re}\int_{\mathbb{R}}y^{2}\Delta_{k}\omega_{k}\overline{\omega_{k}}\mathrm{d}y$
$=-8\mathrm{Re}\langle iky\psi_{k},\partial_{y}\psi_{k}\rangle-4\nu\mathrm{Re}\int_{\mathbb{R}}y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y-2\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}.$
@@
$\int_{\mathbb{R}}y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y=-\int_{\mathbb{R}}(\overline{\omega_{k}}+y\partial_{y}\overline{\omega_{k}})\omega_{k}\mathrm{d}y=-\|\omega_{k}\|_{2}^{2}-\int_{\mathbb{R}}y\partial_{y}\overline{\omega_{k}}\omega_{k}\mathrm{d}y,$
@@
$\mathrm{Re}\int_{\mathbb{R}}y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y=\frac{1}{2}\left(\int_{\mathbb{R}}y\partial_{y}\omega_{k}\overline{\omega_{k}}\mathrm{d}y+\int_{\mathbb{R}}y\partial_{y}\overline{\omega_{k}}\omega_{k}\mathrm{d}y\right)=-\frac{1}{2}\|\omega_{k}\|_{2}^{2}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla_{k}\psi_{k}\|_{2}^{2}=-2\mathrm{Re}\int_{\mathbb{R}}\partial_{t}\Delta_{k}\psi_{k}\overline{\psi_{k}}\mathrm{d}y=-2\mathrm{Re}\int_{\mathbb{R}}(-iky^{2}\omega_{k}+2ik\psi_{k}+\nu\Delta_{k}\omega_{k})\overline{\psi_{k}}\mathrm{d}y=-2\nu\|\omega_{k}\|_{2}^{2}+2\mathrm{Re}\int_{\mathbb{R}}iky^{2}\Delta_{k}\psi_{k}\overline{\psi_{k}}\mathrm{d}y,$
@@
$\mathrm{Re}\int_{\mathbb{R}}iky^{2}\Delta_{k}\psi_{k}\overline{\psi_{k}}\mathrm{d}y=\mathrm{Re}\int_{\mathbb{R}}-\partial_{y}(iky^{2}\overline{\psi_{k}})\partial_{y}\psi_{k}-ik^{3}y^{2}\psi_{k}\overline{\psi_{k}}\mathrm{d}y=2\mathrm{Re}\langle iky\psi_{k},\partial_{y}\psi_{k}\rangle.\qed$
@@
$\frac{\mathrm{d}E_{k}}{\mathrm{d}t}=$ $-\left(c_{\gamma}\gamma_{k}\nu\|\omega_{k}\|_{2}^{2}+\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}+c_{\alpha}\alpha_{k}\nu\|\Delta_{k}\omega_{k}\|_{2}^{2}+4c_{\beta}\beta_{k}|k|^{2}\|y\omega_{k}\|_{2}^{2}+c_{\gamma}\gamma_{k}\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}+8c_{\beta}\beta_{k}|k|^{2}\|\partial_{y}\psi_{k}\|_{2}^{2}\right)$
$-2c_{\alpha}\alpha_{k}\mathrm{Re}\langle iky\omega_{k},\partial_{y}\omega_{k}\rangle-4c_{\beta}\beta_{k}\nu\mathrm{Re}\langle\Delta_{k}\omega_{k},iky\partial_{y}\omega_{k}\rangle.$
@@
$\lambda_{k}E_{k}$ $\lesssim\nu\|k\omega_{k}\|_{2}^{2}+\alpha_{k}\nu\|k\nabla_{k}\omega_{k}\|_{2}^{2}+\gamma_{k}\nu|k|^{2}\|y\omega_{k}\|_{2}^{2}+\gamma_{k}\nu\|k\partial_{y}\psi_{k}\|_{2}^{2}$
$\lesssim\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}+\alpha_{k}\nu\|\Delta_{k}\omega_{k}\|_{2}^{2}+\nu^{\frac{1}{3}}|k|^{2}\|y\omega_{k}\|_{2}^{2}+\gamma_{k}\nu\|\omega_{k}\|_{2}^{2}$
$\lesssim I.\qed$